OpenStudy (anonymous):

How do i solve 3equals 4-7

OpenStudy (immanuelv):

3 = 4-7?

OpenStudy (anonymous):

3=4-7 x=-3+-3 x=-9

OpenStudy (immanuelv):

idk where you found x here.... but okay :/